# ViEEG: Hierarchical Neural Coding with Cross-Modal Progressive Enhancement for EEG-Based Visual Decoding

### v2

@@ Appendix C Whole Results of ablation Study
C.2. ablation on Cross Attention
We further analyzed the effectiveness of the cross-attention module, which is essential for integrating the hierarchical features in ViEEG. Removing the cross-attention component (w/o C-Att) resulted in a noticeable performance drop. In the subject-dependent experiments, the configuration without cross-attention achieved a Top-1 accuracy of 32.4% and Top-5 accuracy of 66.8%, compared to 34.1% and 71.3% for the full ViEEG model. In the subject-independent setting, similar trends were observed, with the full model outperforming the variant without cross-attention by significant margins. This highlights the crucial role of cross-attention in aligning and integrating hierarchical EEG embeddings, thereby improving the robustness and generalization of the model across different subjects.
@@
In Figure 4, we present line charts that highlight the changes in accuracy before and after the integration of cross-attention for each feature. The four subplots represent the Top-1 and Top-5 accuracy changes for each feature: $F_{b}$, $F_{f}$, $F_{r}$, and $F_{EEG}$. These subplots illustrate the performance improvements across the subject-dependent and subject-independent settings. The line charts clearly show that, although individual embeddings benefit from cross-attention, the most significant boost is observed when combining all three embeddings. This visualization allows for a clearer comparison of how cross-attention enhances the overall performance of the ViEEG model across various feature combinations. More detailed results are available in the Appendix for further reference.
Figure 9. Detailed Ablation Study of Attention on Hierarchical Visual Embeddings in ViEEG.
C.3. Parameter Analysis of Attention Modules
Figure 10. Parameter Analysis of Attention Modules in ViEEG. Top row: Performance trends (Top-1, Top-3, Top-5 accuracies) for varying numbers of attention layers under subject-dependent and subject-independent settings. Bottom row: Performance trends for different numbers of attention heads under the same settings.
@@ C.3. Parameter Analysis of Attention Modules
Figure 11. Representational Similarity Matrices Across Six Subjects: ViEEG vs. NICE vs. ATM (Top: ViEEG; Middle: NICE; Bottom: ATM; CB denotes Color Bar).
Appendix D Model Computing Resource Comparison
Methods FLOPs. Param. Avg Top-1 / Top-5 (%)
NICE 42.7354 M 2630.940 K 27.3 / 59.7
ATM 98.209 M 3072.662 K 26.1 / 58.1
CopCap 178.973 M 2954.073 K 33.3 / 60.5
ViEEG 127.973 M 7924.488 K 40.9 / 74.5
Table 7. Comparison of model parameters, computational cost, and performance across different methods.
Table 7 presents a comprehensive comparison of different models in terms of computational cost (FLOPs), parameter count, and recognition performance. Among the baseline methods, CopCap achieves relatively better Top-1/Top-5 accuracy (33.3% / 60.5%) but incurs the highest computational cost (178.973M FLOPs). In contrast, our proposed ViEEG achieves a significant performance gain, reaching 40.9% Top-1 and 74.5% Top-5 accuracy, outperforming all baseline methods by a large margin.
Although ViEEG has a higher parameter count (7924.488K), this increase is attributed to our self-contained hierarchical EEG encoder, which consists of spatiotemporal convolution and cross-attention modules at multiple levels. This design enables more effective extraction and integration of multi-scale EEG representations, which is crucial for capturing the complex brain responses associated with visual stimuli. Despite the higher parameter count, ViEEG maintains a reasonable computational cost (127.973M FLOPs), demonstrating a favorable trade-off between efficiency and accuracy.
Appendix E Representational Analysis
To further support the analysis in the main text, we visualize the representational similarity matrices (RSMs) of ViEEG, NICE (Song et al., 2024), and ATM (Li et al., 2024) across the first six subjects from the THINGS-EEG dataset, as shown in Figure 11. Each matrix reflects how well different EEG conditions are organized in the learned representation space.
@@ Appendix E Representational Analysis
Appendix F Image Retrieval Results of ViEEG
@@
To provide further insight into the effectiveness of our multi-embedding retrieval strategy, we visualize the zero-shot image retrieval results for 15 test samples from Subject 8 in the THINGS-EEG dataset. For each sample, we present the Top-10 retrieval results using four types of EEG-to-image embeddings: binary object mask (BOM), foreground object (FO), raw scene (RS), and their concatenated form (Triple embedding), as shown in Figure 12, Figure 13, Figure 14, and Figure 15, respectively.
@@
Figure 12. Zero-shot Image Retrieval Results Using Binary Object Mask (BOM) Embeddings (Visualizations of the Top-10 retrievals for 8th Subject using binary contour-based EEG-to-image features.)
@@
Figure 13. Zero-shot Image Retrieval Results Using Foreground Object (FO) Embeddings (Visualizations of the Top-10 retrievals for 8th Subject using EEG features aligned with foreground object representations.)
@@
Figure 14. Zero-shot Image Retrieval Results Using Raw Scene (RS) Embeddings (Visualizations of the Top-10 retrievals for 8th Subject based on full-scene EEG-to-image alignment.)
@@
Figure 15. Zero-shot Image Retrieval Results Using Triple Embedding Fusion (BOM + FO + RS) (Visualizations of the Top-10 retrievals for 8th Subject using concatenated multi-scale EEG embeddings.)
Figure 16. Reconstructed images of animal items using ViEEG.)
Figure 17. Reconstructed images of food items using ViEEG.)
Figure 18. Reconstructed images of vehicle items using ViEEG.)
@@
The retrieval results reveal that each embedding modality has its own strengths under different conditions. BOM performs best when the object has simple and distinctive contours (e.g., balls, cones), where boundary information alone is sufficient for recognition. FO excels when the object is visually entangled with the background—such as small or embedded items—where isolating the foreground is critical for accurate decoding. RS generally offers robust performance across diverse scenarios, particularly when the scene-level semantics are dominant and the image embedding is distinctive.
@@
Appendix G Image Reconstruction Results of ViEEG
G.1. Stable Diffusion XL and IP-Adapter
To enable high-quality image reconstruction from EEG signals, we leverage Stable Diffusion XL (SDXL)(Podell et al., 2023), a state-of-the-art generative model known for its ability to produce photorealistic and semantically rich images. To bridge the modality gap between neural signals and visual content, we incorporate the IP-Adapter(Ye et al., 2023), which employs dual cross-attention modules to inject conditioning information into the denoising process. Specifically, we use CLIP-ViT-H/14 (Radford et al., 2021) to encode EEG-derived representations, enabling them to effectively influence the denoising path within the SDXL U-Net, thus guiding the model toward more precise and semantically coherent outputs.
@@ G.1. Stable Diffusion XL and IP-Adapter
G.2. Additional Visualization of ViEEG’s Image Reconstruction
We present additional image reconstruction results from ViEEG on the THINGS-EEG dataset, as shown in Figure 18, Figure 17, and Figure 16. The figures illustrate the reconstruction results in three semantic categories: vehicles, food, and animals.
Overall, ViEEG successfully reconstructs images that maintain semantic consistency with the original stimuli. For example, in Figure 18 (vehicles), reconstructed images effectively capture key structural and shape features of cars, airplanes, and boats. Similarly, Figure 17 (food items) shows that ViEEG can generate food-related textures and shapes, demonstrating its ability to retain category-level information. In Figure 16 (animals), the model reconstructs recognizable features such as fur textures and body structures, reflecting a strong alignment with the original stimuli.
However, certain semantic inconsistencies are observed in some cases. For instance, in Figure 16, when the original image is a cat, some reconstructions depict human faces. This could be due to EEG-induced associative activations, where the subject subconsciously links cats to human interactions. Additionally, errors may arise from attention fluctuations during EEG recording, leading to unintended noise in brain activity. These cases highlight potential challenges in EEG-based visual decoding, emphasizing the need for further refinement in feature alignment and reconstruction fidelity.

### v3

@@ 9.1 Ablation on Hierarchical Embedding
Subject dependent - train and test on one subject
LSTM (Hochreiter 1997) 25.8 60.5 23.1 49.3 30.5 60.8 29.6 60.1 20.3 41.2 25.4 51.7 28.5 59.0 30.6 62.0 27.0 53.1 34.8 67.7 27.6 56.5
DGCNN (Song et al. 2018) 18.2 49.1 14.8 43.2 20.5 52.4 29.0 63.2 12.8 38.8 17.8 53.9 20.6 55.4 29.6 65.4 21.7 52.2 26.1 62.4 21.1 53.6
EEGNet (Lawhern et al. 2018) 21.6 55.4 25.3 59.2 25.3 61.2 30.5 67.1 16.4 46.5 26.7 61.3 25.2 58.6 37.1 71.1 26.0 61.2 32.8 67.6 26.7 60.9
NICE (Song et al. 2024) 31.1 66.7 34.2 65.8 38.9 71.2 45.8 79.5 26.7 69.1 40.6 74.3 35.4 70.6 50.2 80.5 35.1 72.3 40.0 75.7 37.8 72.6
ATM (Li et al. 2024) 30.8 67.3 34.8 65.6 36.6 72.3 44.9 80.6 28.2 60.9 37.5 72.2 39.1 74.6 49.3 80.9 34.2 71.6 40.7 77.8 37.6 72.4
CogCap (Zhang et al. 2024) 30.2 65.6 33.1 64.1 37.4 70.8 43.8 78.9 25.7 57.4 34.6 63.5 35.9 71.1 48.7 78.8 35.6 70.3 38.2 76.5 36.3 69.7
Subject independent - leave one subject out for test
LSTM (Hochreiter 1997) 14.8 38.1 13.8 32.5 8.0 24.6 13.0 35.8 10.2 27.6 12.2 32.6 8.5 26.5 14.3 38.8 6.5 28.2 14.1 37.4 11.5 32.2
DGCNN (Song et al. 2018) 12.9 36.5 17.5 39.3 6.3 25.5 12.8 31.0 9.2 32.0 13.7 36.9 7.3 28.8 14.3 32.8 14.2 36.4 15.6 41.3 12.4 34.1
EEGNet (Lawhern et al. 2018) 13.5 36.2 15.9 40.7 7.6 27.8 16.0 37.4 10.3 30.0 12.2 38.2 8.9 29.5 16.8 39.5 10.5 30.6 21.2 51.0 13.3 36.1
NICE (Song et al. 2024) 18.4 50.3 21.9 47.3 14.5 44.6 22.1 53.9 15.0 39.8 20.2 48.4 16.3 43.1 18.4 44.6 19.7 45.2 27.8 45.5 19.4 46.3
ATM (Li et al. 2024) 20.7 48.0 22.2 49.6 14.3 39.6 20.5 48.2 18.2 43.0 18.0 44.1 20.5 42.3 21.6 46.3 19.1 41.6 27.2 55.7 19.8 45.8
CogCap (Zhang et al. 2024) 19.1 47.4 17.1 44.5 10.2 29.0 20.3 44.2 11.2 32.9 17.3 38.6 15.3 35.5 17.4 39.7 13.8 38.8 22.5 50.6 16.4 40.1
Table 7: Ablation study of EEG encoder: Top-1/Top-5 object recognition accuracy(%).
Subject dependent - train and test on one subject
[…]
Subject independent - leave one subject out for test
[…]
Table 8: Ablation study of hierarchical processing order: Top-1/Top-5 object recognition accuracy(%).
9.2 Ablation on Cross Attention
We further analyzed the effectiveness of the cross-attention module, which is essential for integrating the hierarchical features in ViEEG. Removing the cross-attention component (w/o C-Att) resulted in a noticeable performance drop. In the subject-dependent experiments, the configuration without cross-attention achieved a Top-1 accuracy of 32.4% and Top-5 accuracy of 66.8%, compared to 34.1% and 71.3% for the full ViEEG. In the subject-independent setting, similar trends were observed, with the full model outperforming the variant without cross-attention by significant margins. This highlights the crucial role of cross-attention in aligning and integrating hierarchical EEG embeddings, thereby improving the robustness and generalization of the model across different subjects.
@@ 9.2 Ablation on Cross Attention
In Figure 4, we present line charts that highlight the changes in accuracy before and after the integration of cross-attention for each feature. The four subplots represent the Top-1 and Top-5 accuracy changes for each feature: $F_{b}$, $F_{f}$, $F_{r}$, and $F_{EEG}$. These subplots illustrate the performance improvements across the subject-dependent and subject-independent settings. The line charts clearly show that, although individual embeddings benefit from cross-attention, the most significant boost is observed when combining all three embeddings. This visualization allows for a clearer comparison of how cross-attention enhances the overall performance of ViEEG across various feature combinations. More detailed results are available in the Appendix for further reference.
9.3 Ablation on EEG Encoder
We conducted a detailed ablation study of EEG Encoder. For ViEEG, the EEG feature encoder plays a crucial role as same as image segmentation, and it is the foundation for further modal alignment. Replacing our EEG encoder modules with other EEG encoders (e.g., ATM (Li et al. 2024) and NICE (Song et al. 2024)) for feature extraction led to an average performance significantly drop, as shown from Table 7 below. Our proposed ViEEG achieves the best performance in every experiment, which demonstrates that the SOTA performance of our method is not only dependent on image segmentation features, but also on the hierarchical EEG Encoder. Moreover, as shown by the “increase” value from tables, our proposed hierarchical visual decoding framework shows a significantly improvement in the accuracy rate of object recognition. Thus, ViEEG’s SOTA performance arises from the synergy between biologically motivated segmentation and hierarchical EEG encoding.
9.4 Ablation on Hierarchical Processing Order
While our framework adopts a bottom-up order (contour $\to$ object $\to$ scene) based on visual neuroscience, we also tested other alternative integration orders (e.g., object $\to$ contour $\to$ scene, random) in the Table 8. Results show that the biologically inspired order consistently outperforms others, suggesting that respecting the natural visual hierarchy better aligns EEG features with semantic structure.
Figure 9: Parameter analysis of attention modules in ViEEG. Top row: Performance trends (Top-1, Top-3, Top-5 accuracies) for varying numbers of attention layers under subject-dependent and subject-independent settings. Bottom row: Performance trends for different numbers of attention heads under the same settings.
9.5 Parameter Analysis of Attention Modules
@@
Figure 10: Representational similarity matrices across first six subjects: ViEEG vs. NICE vs. ATM (CB: Color bar).
@@
10 Representational Analysis
To further support the analysis in the main text, we visualize the representational similarity matrices (RSMs) of ViEEG, NICE (Song et al. 2024), and ATM (Li et al. 2024) across the first six subjects from the THINGS-EEG dataset, as shown in Figure 10. Each matrix reflects how well different EEG conditions are organized in the learned representation space.
@@ 10 Representational Analysis
11 Model Computing Resource Comparison
Methods FLOPs. Param. Train time (s) Test time (s)
NICE 42.7354 M 2630.940 K 0.0412 0.0058
ATM 98.209 M 3072.662 K 0.0524 0.0088
CopCap 178.973 M 2954.073 K 0.0591 0.0056
ViEEG 127.973 M 7924.488 K 0.0682 0.0062
Table 9: Comprehensive comparison of computational requirements and efficiency.
[…]
NICE (Song et al. 2024) 0.353 0.774 0.87 0.736 0.789 0.825 0.588
ATM (Li et al. 2024) 0.351 0.779 0.873 0.745 0.792 0.823 0.591
CogCap (Zhang et al. 2024) 0.346 0.768 0.866 0.737 0.781 0.832 0.595
ViEEG 0.356 0.787 0.881 0.748 0.798 0.811 0.584
Table 10: Quantitative assessments of reconstruction quality.
Table 9 presents an integrated analysis of computational requirements and recognition performance. The comparison reveals three key insights through the following aspects:
Computational Efficiency and Accuracy Balance: CopCap achieves the highest baseline accuracy (33.3% Top-1/60.5% Top-5) but requires significantly more computations than other methods, with 179.0M FLOPs – over three times that of NICE (42.7M FLOPs, 27.3% Top-1). Our ViEEG demonstrates a balanced design: while using 28% fewer computations than CopCap (128.0M vs 179.0M FLOPs), it achieves a remarkable 40.9% Top-1 accuracy (7.6% improvement). Notably, ViEEG maintains competitive inference speed (6.2ms), being only marginally slower than CopCap’s 5.6ms despite substantial accuracy gains.
Parameter Utilization Effectiveness: The parameter counts reflect distinct design strategies: NICE (2.63M parameters) prioritizes lightweight design but limits performance (27.3% Top-1). CopCap (2.95M parameters) slightly increases parameters to improve temporal modeling (33.3% Top-1). ViEEG (7.92M parameters) strategically allocates parameters to hierarchical components, enabling multi-scale EEG feature integration. This design yields 74.5% Top-5 accuracy (+14% over CopCap), validating our spatial-frequency interaction mechanism.
Practical Deployment Considerations: ViEEG maintains practical efficiency across metrics: Training time per step (68.2ms) shows only moderate increase compared to CopCap (59.1ms). Inference latency (6.2ms) closely matches fastest baselines (NICE:5.8ms, CopCap:5.6ms). Accuracy superiority: Achieves 13.6% Top-1 improvement over NICE with negligible 0.4ms inference delay. The results validate that ViEEG successfully balances model capacity and efficiency through its hierarchical architecture, delivering state-of-the-art performance while maintaining practical deployment feasibility.
The results demonstrate ViEEG’s ability to convert increased parameters into significant accuracy enhancements (40.9% vs 33.3% Top-1) while avoiding excessive computational costs. Compared to CopCap’s 179.0M FLOPs, ViEEG reduces computations by 51.0M FLOPs (equivalent to 28.5% reduction) while improving Top-1 accuracy by 7.6%, showing superior efficiency-accuracy trade-off. The test time comparison further confirms that ViEEG’s accuracy gains (74.5% vs 60.5% Top-5) come without sacrificing real-time processing capability, making it suitable for practical brain signal analysis applications.
12 Image Retrieval Results of ViEEG
@@
To provide further insight into the effectiveness of our multi-embedding retrieval strategy, we visualize the zero-shot image retrieval results from Subject 8 in the THINGS-EEG dataset. For each sample, we present the Top-10 retrieval results using four visual embeddings: binary object mask (BOM), foreground object (FO), raw scene (RS), and their concatenated form (Triple embedding), as shown in Figure 11, Figure 12, Figure 13, and Figure 14, respectively.
@@
Figure 11: Zero-shot image retrieval results for 15 test samples using binary object mask (BOM) embeddings (Visualizations of the Top-10 retrievals for Subject-8 using binary contour-based EEG-to-image features).
@@
Figure 12: Zero-shot image retrieval results for 15 test samples using foreground object (FO) embeddings (Visualizations of the Top-10 retrievals for Subject-8 using EEG features aligned with foreground object representations).
@@
Figure 13: Zero-shot image retrieval results for 15 test samples using raw scene (RS) embeddings (Visualizations of the Top-10 retrievals for Subject-8 based on full-scene EEG-to-image alignment).
@@
Figure 14: Zero-shot image retrieval results for 15 test samples using triple embedding fusion (BOM + FO + RS) (Visualizations of the Top-10 retrievals for Subject-8 using concatenated multi-scale EEG embeddings).
Figure 15: Reconstructed images of animals items using ViEEG.
Figure 16: Reconstructed images of foods items using ViEEG.
Figure 17: Reconstructed images of vehicles items using ViEEG.
@@
The retrieval results reveal that each embedding modality has its own strengths under different conditions. BOM performs best when the object has simple and distinctive contours (e.g., balls, cones), where boundary information alone is sufficient for recognition. FO excels when object is visually entangled with the background—such as small or embedded items—where isolating the foreground is critical for accurate decoding. RS generally offers robust performance across diverse scenarios, particularly when scene-level semantics are dominant and image embedding is distinctive.
@@
13 Image Reconstruction Results of ViEEG
13.1 Stable Diffusion XL and IP-Adapter
To enable high-quality image reconstruction from EEG, we leverage Stable Diffusion XL (SDXL) (Podell et al. 2023), a state-of-the-art generative model known for its ability to produce photorealistic and semantically rich images. To bridge the modality gap between neural signals and visual content, we incorporate the IP-Adapter (Ye et al. 2023), which employs dual cross-attention modules to inject conditioning information into the denoising process. Specifically, we use CLIP-ViT-H/14777https://huggingface.co/laion/CLIP-ViT-H-14-laion2B-s32B-b79K (Radford et al. 2021) to encode image representations, enabling them to effectively influence the denoising path within the SDXL, thus guiding the model toward more precise and semantically coherent outputs.
@@ 13.1 Stable Diffusion XL and IP-Adapter
13.2 Image Reconstruction Results of ViEEG
We present additional image reconstruction results from ViEEG on the THINGS-EEG dataset, as shown in Figure 17, Figure 16, and Figure 15. The figures illustrate the reconstruction results in three semantic categories: vehicles, food, and animals. Overall, ViEEG successfully reconstructs images that maintain semantic consistency with the original stimuli. For example, in Figure 17 (vehicles), reconstructed images effectively capture key structural and shape features of cars, airplanes, and boats. Similarly, Figure 16 (food items) shows that ViEEG can generate food-related textures and shapes, demonstrating its ability to retain category-level information. In Figure 15 (animals), the model reconstructs recognizable features such as fur textures and body structures, reflecting a strong alignment with the original stimuli.
However, certain semantic inconsistencies are observed in some cases. For instance, in Figure 15, when the original image is a cat, some reconstructions depict human faces. This could be due to EEG-induced associative activations, where the subject subconsciously links cats to human interactions. Additionally, errors may arise from attention fluctuations during EEG recording, leading to unintended noise in brain activity. These cases highlight potential challenges in EEG visual decoding, emphasizing the need for further refinement in feature alignment and reconstruction fidelity.
13.3 Quantitative Metrics of Reconstructed Image
We now provide quantitative evaluation of EEG-to-image reconstruction in Table 10, including standard metrics used in prior work (e.g., ATM (Li et al. 2024)). These include:
  • SSIM: measures low-level structural similarity (e.g., texture, edges);
  • AlexNet (2/5) and Inception: assess mid/high-level perceptual feature similarity;
  • CLIP: evaluates alignment in joint vision-language space, relevant for semantic fidelity;
  • EffNet and SwAV: capture feature compactness and unsupervised visual consistency.
Although the same diffusion model is used, ViEEG achieves consistently better scores due to improved EEG-CLIP alignment. This suggests that our hierarchical EEG encoder contributes to more semantically faithful reconstructions.